OpenStudy (anonymous):

Solve using the substitution method. x-2y=5 2x-y=16

OpenStudy (anonymous):

Solve both equations for x: x = 5 + 2y x = (16 + y)/2 Substitute the x for the other equation and solve for y: 5 + 2y = (16 + y)/2 5 + 2y = 8 + y/2 2y = 3 + y/2 3y/2 = 3 3y = 6 y = 2

OpenStudy (anonymous):

using the ist eq. x-2y=5 x=2y+5 substitute this to the 2nd eq 2x-y=16 that is 2(2y+5)-y=16 4y+10-y =16 3y+10=16 3y=16-10 3y=6 y=6/3 y=2 now subs his to the ist eq or 2nd eq,, im choosing the ist eq. x-2y=5 therefore x-2(2)=5 x-4 = 5 x=5+4=9 answ..